OpenStudy (anonymous):

1. explain which measure off central tendency is most affected by an extreme value. 2.write a set of data with at least four numbers that has a mean of 8 and a median is not 8. find the mean median and mode round if necessary. 3. 4,5,7,3,9,11,23,37 4. 7.2,3.6,9.0,5.2,7.2,6.5,3.6

OpenStudy (anonymous):

1. mean 2. 2 4 10 10 14 3. mean = 99/8, median = (7+9) / 2 = 8, no mode